OpenStudy (anonymous):

about 11/20 of students ages 10 to 17 name math, science, or art as their favorite subject in school. Art is the favorite subject for about 3/50 of the students ages 10 to 17. for what fractin of students this age is math or schience their favorite subject?

OpenStudy (mertsj):

\[\frac{11}{20}-\frac{3}{50}=\frac{49}{100}\]

OpenStudy (anonymous):

That was correct! Awsome thanks.

OpenStudy (anonymous):

Stick around you might come in handy bro!